OpenStudy (anonymous):

determine the slope of a line passing through the points (12, 5) and (6, 7)

OpenStudy (btaylor):

\[slope = \frac{\Delta y}{\Delta x}= \frac{y_2-y_1}{x_2-x_1} = \frac{7-5}{6-12} = \frac{2}{-6} = - \frac{1}{3}\]